OpenStudy (anonymous):

The coefficient of x^2 in the expansion (k+1/3x)^5 is 30. Find the value of the constant k. I have almost absolutely no idea how to do this... Please help!

OpenStudy (anonymous):

This involves pascall's triangle.

OpenStudy (anonymous):

What exactly about it is Pascal, is it something to do with binomial expansion?

OpenStudy (anonymous):

\[ Ck^3\left(\frac 13 x\right)^2 = 30x^2 \]

OpenStudy (anonymous):

In this case \[ C = \binom 53=\binom 52 \]

OpenStudy (anonymous):

You could use Pascal's triangle to find the binomial coefficient.

OpenStudy (anonymous):

Or you could use the combination function if you already know it.

OpenStudy (anonymous):

So would it be something like 5C3 x k^2 x (1/3x)^2=30?

OpenStudy (anonymous):

=30x^2

OpenStudy (anonymous):

Why is it 30x^2?

OpenStudy (anonymous):

Trust me, you don't want \(x\) in your answer. It said \(30\) is only the coefficient of the \(x^2\) term, so the whole term is \(30x^2\)

OpenStudy (anonymous):

Hmm ok, so 5C3 x k^2 x (1/3x)^2=30x^2 Do I just solve that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok thank you for your help.

OpenStudy (anonymous):

Thanks so much!!!